# Black Hole Entropy, Quantum Corrections and EFT Transitions

### v1

@@ 3.2.2 The transition regime
The analysis presented in Sections 3.1 and 3.2.1 is very instructive and teaches us that, in those cases where our description can be reliably embedded into the large volume patch, the four-dimensional EFT provides sensible answers for the relevant physical properties of the supersymmetric black hole solutions considered therein. However, we noticed that the additional contributions arising from higher-derivative terms included in the (half-)superspace integral (2.5) lead to a certain tower of 4d local operators involving the (anti-sefl-dual part of the) graviton and graviphoton field strengths, which correct all these quantities via some perturbative series depending on a real-valued expansion parameter $\alpha=1/Y^{0}$, cf. eq. (3.16). Our main concern in what follows will be to understand both its physical significance as well as whether one could reach some pathological regime where the series would seem to break down, thus invalidating the four-dimensional effective description of the BPS system.
@@
where $r_{5}$ is the physical radius of the dual M-theory circle evaluated at the horizon —as computed from the D0 mass, and $r_{h}$ denotes that of the black hole. Note that the last identity in (3.23) readily follows from the identification $r_{h}=|\mathscr{Z}|\kappa_{4}/\sqrt{8\pi}$ (cf. eq. (2.19)), as well as the fact that the internal $\mathbf{S}^{1}$ radius is captured by the inverse mass of the Kaluza-Klein replica, which in the present context corresponds to the D0-brane states. The latter have a (running) mass that is easily determined to be $m_{\rm D0}=\sqrt{8\pi}|X^{0}|e^{\mathscr{K}/2}/\kappa_{4}$. Hence, when evaluated at the attractor point, the (absolute value of the) expansion parameter precisely captures the relative size between the black hole horizon and M-theory circle.
@@
Figure 1: Schematic depiction of the spatial profile for the M-theory circle (blue) in the 4d supersymmetric black hole background. Its asymptotic size is determined by the vacuum expectation value of the Calabi–Yau volume (in string units), and it evolves smoothly according to the attractor flow towards its fixed value at the horizon. If the size of the black hole is comparable to that of the extra circle, KK gravitons can be easily excited, yielding large corrections to e.g., the entropy.
Therefore, recall from our discussion in Section 3.2.1 that the regime where the 4d EFT seemed to organize itself in a perturbative and well-behaved way so as to correctly reproduce the physical properties of the D0-D2-D4 black hole system, occurred precisely when $\alpha\ll 1$. When this is the case, the black hole becomes much bigger than the size of the extra dimension, such that a four-dimensional effective description must be able to describe the relevant physics. On the other hand, if we now consider the opposite situation where both radii are close to each other —namely when $r_{5}\gtrsim r_{h}$, then a putative 5d picture seems to be required. This corresponds to black hole solutions with $\alpha\gtrsim\mathcal{O}(1)$ at the horizon, and for those something interesting must be going on, since the series controlling $G(\alpha,\Upsilon)$ breaks down very quickly. This stems from the fact that the perturbative expansion capturing the quantum deformations of the black hole solutions exhibit numerical coefficients that grow in a factorial way, namely $c_{g-1}^{3}\sim(2g-3)!$ Consequently, the series expansion
@@
All in all, we conclude that the point $|\alpha|\gtrsim\mathcal{O}(1)$ marks some sharp transition Calderon-Infante:2025ldq where the 4d EFT —including the tower of higher-curvature and higher-derivative local operators derived from (2.5)— provides unphysical results, at least with regard to certain black hole properties. However, the interpretation of this failure is perfectly reasonable according to our discussion above. Indeed, the four-dimensional supergravity theory starts giving misleading predictions for the physics associated to certain BPS black holes precisely when these attain sizes which are of the order of the extra compact dimension (in the dual M-theory picture). At this point, one can no longer view the internal circle to be adiabatically fibered over the spatial $\mathbb{R}^{3}$ external to the horizon (see Figure 1), and in fact local fluctuations in the black hole geometry can easily excite KK modes (i.e., D0-branes). Hence, one should not expect a purely 4d description to be able to capture the physical properties associated to these systems, since they already belong to the five-dimensional realm. On the other hand, the higher-dimensional EFT should be able to cure somehow this pathology upon including highly-non local effects (when seen from the 4d perspective), which is what we will devote all our efforts to in the upcoming sections.
@@ 3.3.1 A resummed entropy formula
where one should substitute the particular value of $Y^{0}$ that solves the attractor equation displayed in (3.11). This expression should be moreover understood as the resummed version of eq. (3.19) above.
Let us also note, in passing, that the present analysis reinforces the idea that the smallest possible black hole size is attained when the linear term in the generalized prepotential (2.31) becomes of the same order as (or even dominates over) the classical cubic piece. This happens whenever the magnetic charges $p^{a}$ are all of order one, and in this case, eq. (3.33) precisely accounts for the minimal black hole entropy. The latter provides an $\mathcal{O}(\Lambda_{\text{QG}}^{2}/M_{\text{Pl,}\,4}^{2})$ number, with $\Lambda_{\text{QG}}$ denoting here the quantum gravity cut-off, i.e., the 5d Planck scale) Cribiori:2022nke ; Calderon-Infante:2023uhz ; smallBHsMunich ; 4dBHs .
@@ 3.3.2 Explicit gluing with the 5d solution
As is well known, the four-dimensional EFT considered so far (cf. Section 2.1) has a natural embedding into a 5d theory with one direction compactified on a circle Witten:1995zh ; Cadavid:1995bk . At the classical —i.e., two-derivative— level, the supersymmetric black hole solutions studied in this section can be uplifted to five-dimensional, supersymmetric black strings wrapping the internal periodic direction Maldacena:1996ky . Therefore, in the limit of infinite compactification radius they would appear to extend indefinitely. In string theory, this scenario is realized by uplifting Type IIA compactified on a Calabi–Yau threefold to M-theory reduced on the same compact space times a circle, and subsequently taking the large $\mathbf{S}^{1}$ limit. These solutions exist classically for every value of the black hole and the (asymptotic) compactification circle radii Gauntlett:2002nw . However, one could naively wonder whether quantum corrections could spoil them. Importantly, notice that these two pictures must be regarded as complementary, limiting descriptions of the same physical object, since they are naturally described using two different EFTs. In particular, the 4d black hole is valid as a four-dimensional EFT solution as long as the quantum corrections are (highly) suppressed, which is equivalent to being in the regime where $\alpha\ll 1$ at the horizon. On the other hand, for $\alpha\gg 1$ the correct EFT description is the one of a five-dimensional black string wrapped on the M-theory circle, with the horizon transverse to and much smaller than the latter. Finally, for $\alpha\sim\mathcal{O}(1)$ the physical object still requires a 5d EFT description where the quantum corrections associated to the compactification circle are not necessarily small and need to be properly taken into account.
@@
where $c_{L}$ is the (left) central charge of the sigma model associated to the moduli space of the worldvolume theory of the branes whose gravitational backreation generates the black hole. The central charge can be evaluated to be Maldacena:1997de ; Vafa:1997gr ; Harvey:1998bx
@@
Interestingly, and in contrast to the four-dimensional case, in five dimensions one can actually prove that the entropy has the structure (3.34) also with a macroscopic computation. We start by clarifying what is the interpretation of the macroscopic entropy of a black string and, in general, of any extended black $p$-brane. For simplicity, we discuss this point in the two-derivative approximation. Clearly, if the entropy were simply the analog of the Bekenstein–Hawking area law we would obtain that all extended black-branes would have infinite entropy. Therefore, the way to obtain meaningful thermodynamics for such objects is to introduce regularized quantities in the form of worldvolume densities (see, for instance, deAntonioMartin:2012bi ; Meessen:2012su ). Reintroducing Newton’s constant, we can conveniently define the entropy density $s$ for a $p$-brane living in $d$ spacetime dimensions as
@@
where $A_{\text{hor}}$ is the $(d-p-2)$-dimensional black-brane transverse horizon, and $G_{d}$ is the $d$-dimensional Newton constant. Hence, for a five-dimensional black string one obtains
@@
where $\mathcal{R}_{\mu\nu}$ is the Ricci tensor. Equation (3.41) is not completely fixed by the near-horizon geometry and to evaluate it we must know the precise structure of $\mathcal{L}_{3}$. Interestingly, this computation correctly reproduces the central charge (3.35) upon considering just the standard, two-derivative 5d $\mathcal{N}=1$ supergravity action supplemented with the known four-derivative corrections Castro:2007sd . Consequently, the macroscopic computation can be regarded as one-loop exact. On top of that, this confirms that (3.36) gives not only the leading term in the decompactification limit, but actually the exact entropy of a five-dimensional black string extended along an infinitely long compact circle of volume equal to $\mathcal{V}_{\rm h}$ in 5d Planck units.
@@ 3.4 Including non-perturbative effects
Our aim in the following will be to reconsider this point and show that, in fact, the main conclusions drawn from last section are left unchanged. To do so, we take two alternative routes that ultimately lead to the same answer. The first one proceeds, as explained in Section 3.4.1, by carefully evaluating the Schwinger determinant —in complex proper time— for each state in the D0-brane tower independently (see also Hattab:2024ewk ; Hattab:2024ssg for an independent, more general derivation). Conversely, in Section 3.4.2, we derive an equivalent prescription by resumming the full tower of one-loop contributions that treats both perturbative and non-perturbative corrections on an equal footing. We will also briefly discuss the limitations of this second approach, which is very reminiscent of the recent proposal for computing the non-perturbative topological string partition function put forward in Hattab:2024ewk ; Hattab:2024chf ; Hattab:2024yol ; Hattab:2024ssg .
@@ 3.4.1 Direct evaluation of the Schwinger integral
Our first strategy to obtain a non-perturbative definition of the leading-order prepotential at large volume proceeds similarly as we did in Section 3.3.1. There, following Gopakumar:1998ii ; Gopakumar:1998jq , we showed explicitly how by resorting to the dual M-theory description one is able to rewrite the relevant asymptotic series in an integral Schwinger-like form, cf. eqs. (3.25) and (3.26). Next, we performed the integration using an appropriate change of variable and a certain Fourier transform, which lead us directly to the resummed perturbative expression (3.29). However, in doing so we were not concerned with some subtleties related to both the state-dependent change of integration variable (cf. footnote 3), as well as to possible singularities that could arise within the complex $s$-plane. In fact, it is easy to realize that, when relabeling $s\to s/2\pi n$ so as to reach the l.h.s. of (3.28), we must separate between states with $n\geq 0$ and $n<0$, since their contours cannot be simply deformed into one another due to the (double) poles arising from the hyperbolic sine in the denominator of the integrand. Hence, taking this into account leads to the following two distinct contributions within $\mathcal{I}(\alpha)$
@@
As a next step, and in order to be able to perform the Poisson resummation that lead to the r.h.s. of (3.28), we need to deform the contour integral (3.42b) within the upper half plane so that it coincides with the positive real axis, thereby implying that we should pick up the residues of the infinitely many poles located $s=\frac{2\pi ik}{\alpha}$ for each $k\in\mathbb{N}$ (see Figure 4). The latter give rise to a non-perturbative contribution of the form [^17] Notice that (3.43) can be rewritten in terms of a single function $\varrho(\alpha)=i\frac{\alpha^{2}}{2\pi}\,\sum_{k=1}^{\infty}k^{-2}\left(1-e^{4\pi^{2}k/\alpha}\right)^{-1}$ as follows $\mathcal{I}^{(np)}(\alpha)=\left(1+\alpha\frac{d}{d\alpha}\right)\varrho(\alpha)\,.$
@@
For illustrative purposes, we have depicted the exact non-perturbative contribution, computed from (3.43), in Figure LABEL:sfig:ImaginarypartIcurve. Notice, in particular, the polynomial dependence with the expansion parameter $\alpha$ that arises in the deep five-dimensional regime, whose physical origin may seem surprising from the perspective of the auxiliary topological string theory. Furthermore, one might object that such divergent behavior could potentially undermine the discussion presented in Section 3.3.2, where it was argued that in the $\alpha\to\infty$ limit the entropy formula should match the one computed directly within the uplifted 5d supergravity theory. Very remarkably, we observe that this additional non-perturbative correction in $G(Y^{0},\Upsilon)$ does not modify the attractor equations nor the entropy function associated to the black hole system we are interested in here, since only the real part of $\mathcal{I}(\alpha)$ contributes to those, cf. eqs. (3.11)-(3.14). On hindsight, we should have indeed expected this to happen, since the objects under study are extremal, BPS black holes and thus must be protected by supersymmetry. In any case, it is interesting to see explicitly how this expectation is borne out in the present set-up, also encouraging us to propose that a similar phenomenon would also occur for other four-dimensional, supersymmetric solutions (see Section 4 for more evidence on this).
Let us note, in passing, that the present analysis is also in agreement with recent results obtained in Lin:2024jug , where it was shown that extremal Reissner-Nördstrom black holes exhibit a non-trivial spatial profile for their decay rate induced by non-perturbative emission of Swchinger pairs due to charged particles already existing in the theory, unless the latter are also extremal. Therefore, given that the both the black hole solutions considered herein and the charged D0-branes fulfill this condition Gendler:2020dfp , it makes perfect sense that such a decay channel does not indeed exist.

### v2

@@ 3.2.2 The transition regime
The analysis presented in Sections 3.1 and 3.2.1 is very instructive and teaches us that, in those cases where our description can be reliably embedded into the large volume patch, the four-dimensional EFT provides sensible answers for the relevant physical properties of the supersymmetric black hole solutions considered therein. However, we noticed that the additional contributions arising from higher-derivative terms included in the half-superspace integral (2.5) lead to a certain tower of 4d local operators involving the (anti-sefl-dual part of the) graviton and graviphoton field strengths, which correct all these quantities via some perturbative series depending on a real-valued expansion parameter $\alpha=1/Y^{0}$, cf. eq. (3.16). Our main concern in what follows will be to understand both its physical significance as well as whether one could reach some pathological regime where the series would seem to break down, thus invalidating the four-dimensional effective description of the BPS system.
@@
where $r_{5}$ is the physical radius of the dual M-theory circle evaluated at the horizon —as computed from the D0 mass, and $r_{h}$ denotes that of the black hole. Note that the last identity in (3.23) readily follows from the identification $r_{h}=|\mathscr{Z}|\kappa_{4}/\sqrt{8\pi}$ (cf. eq. (2.19)), as well as the fact that the internal $\mathbf{S}^{1}$ radius is captured by the characteristic Compton wavelength of the Kaluza-Klein replica, which in the present context correspond to the D0-brane states. The latter have a (running) mass that is easily determined to be $m_{\rm D0}=\sqrt{8\pi}|X^{0}|e^{\mathscr{K}/2}/\kappa_{4}$. Hence, when evaluated at the attractor point, the (absolute value of the) expansion parameter precisely captures the relative size between the black hole horizon and M-theory circle.
@@
Figure 1: Schematic depiction of the spatial profile for the M-theory circle (blue) in the 4d supersymmetric black hole background. Its asymptotic size is determined by the vacuum expectation value of the Calabi–Yau volume (in string units), and it evolves smoothly according to the attractor flow towards its fixed value at the horizon. If the size of the black hole is comparable to that of the extra circle, $K$ gravitons can be easily excited, yielding large corrections to e.g., the entropy.
Therefore, recall from our discussion in Section 3.2.1 that the regime where the 4d EFT seemed to organize itself in a perturbative and well-behaved way so as to correctly reproduce the physical properties of the D0-D2-D4 black hole system, occurred precisely when $\alpha\ll 1$. When this is the case, the black hole becomes much bigger than the size of the extra dimension, such that a four-dimensional effective description must be able to describe the relevant physics. On the other hand, if we now consider the opposite situation where both radii are close to each other —namely when $r_{5}\gtrsim r_{h}$, then a putative 5d picture seems to be required. This corresponds to black hole solutions with $\alpha\gtrsim\mathcal{O}(1)$ at the horizon, and for those something interesting must be going on, since the series controlling $G(\alpha,\Upsilon)$ breaks down very quickly. This stems from the fact that the perturbative expansion capturing the quantum deformations of the black hole solutions exhibit numerical coefficients that grow in a factorial way, namely $c_{g-1}^{3}\sim(2g-3)!$ . Consequently, the series expansion
@@
All in all, we conclude that the point $|\alpha|=\mathcal{O}(1)$ marks some sharp transition Calderon-Infante:2025ldq where the 4d EFT —including the tower of higher-curvature and higher-derivative local operators derived from (2.5)— provides unphysical results, at least with regard to certain black hole properties. However, the interpretation of this failure is perfectly reasonable according to our discussion above. Indeed, the four-dimensional supergravity theory starts giving misleading predictions for the physics associated to certain BPS black holes precisely when these attain sizes which are of the order of the extra compact dimension (in the dual M-theory picture). At this point, one can no longer view the internal circle to be adiabatically fibered over the spatial $\mathbb{R}^{3}$ external to the horizon (see Figure 1), and in fact local fluctuations in the black hole geometry can easily excite KK modes (i.e., D0-branes). Hence, one should not expect a purely 4d description to be able to capture the physical properties associated to these systems, since they already belong to the five-dimensional realm. On the other hand, the higher-dimensional EFT should be able to cure somehow this pathology upon including highly-non local effects (when seen from the 4d perspective), which is what we will devote all our efforts to in the upcoming sections.
@@ 3.3.1 A resummed entropy formula
where one must substitute the particular value of $Y^{0}$ that solves the attractor equation displayed in (3.11). This expression should be moreover understood as the resummed version of eq. (3.19) above.
Let us also note, in passing, that the present analysis reinforces the idea that the smallest possible black hole size is attained when the linear term in the generalized prepotential (2.31) becomes of the same order as (or even dominates over) the classical cubic piece. This happens whenever the magnetic charges $p^{a}$ are all of order one, and in this case, eq. (3.33) precisely accounts for the minimal black hole entropy. The latter provides an $\mathcal{O}(M_{\text{Pl,}\,4}^{2}/\Lambda_{\text{QG}}^{2})$ number, with $\Lambda_{\text{QG}}$ denoting here the quantum gravity cut-off, i.e., the 5d Planck scale) Cribiori:2022nke ; Calderon-Infante:2023uhz ; Calderon-Infante:2025pls ; 4dBHs .
@@ 3.3.2 Explicit gluing with the 5d solution
As is well known, the four-dimensional EFT considered so far (cf. Section 2.1) has a natural embedding into a 5d theory with one direction compactified on a circle Witten:1995zh ; Cadavid:1995bk . At the classical —i.e., two-derivative— level, the supersymmetric black hole solutions studied in this section can be uplifted to five-dimensional, supersymmetric black strings wrapping the internal periodic direction Maldacena:1996ky . Therefore, in the limit of infinite compactification radius they would appear to extend indefinitely. In string theory, this scenario is realized by uplifting Type IIA compactified on a Calabi–Yau threefold to M-theory reduced on the same compact space times a circle, and subsequently taking the large $\mathbf{S}^{1}$ limit. These solutions exist classically for every value of the black hole and the (asymptotic) compactification circle radii Gauntlett:2002nw . However, one could naively wonder whether quantum corrections could spoil them. Importantly, notice that these two pictures must be regarded as complementary, limiting descriptions of the same physical object, since they naturally arise upon using two different EFTs. In particular, the 4d black hole is valid as a four-dimensional EFT solution as long as the quantum corrections are (highly) suppressed, which is equivalent to being in the regime where $\alpha\ll 1$ at the horizon. On the other hand, for $\alpha\gg 1$ the correct EFT description is the one of a five-dimensional black string wrapped on the M-theory circle, with the horizon transverse to and much smaller than the latter. Finally, for $\alpha=\mathcal{O}(1)$ the physical object still requires a 5d EFT description where the quantum corrections associated to the compactification circle are not necessarily small and need to be properly taken into account.
@@
where $c_{L}$ is the (left) central charge [^18] For unitary conformal theories, the central charge must satisfy $c_{L}\geq 0$ Cappelli:1986hf ; DiFrancesco:1997nk . From the present, geometrical perspective, this is ensured by the fact that (3.35) equals the second Chern number of the 4-cycle class $P\subset X_{3}$ wrapped by the M5-branes, which is non-negative for nef divisors Miyaoka1987 ; kanazawa2013 of the sigma model associated to the moduli space of the worldvolume theory of the branes whose gravitational backreation generates the black hole. The central charge can be evaluated to be Maldacena:1997de ; Vafa:1997gr ; Harvey:1998bx
@@
Interestingly, and in contrast to the four-dimensional case, in five dimensions one can actually prove that the entropy has the structure (3.34) also with a macroscopic computation. We start by clarifying what is the interpretation of the macroscopic entropy of a black string and, in general, of any extended black $p$-brane. For simplicity, we discuss this point in the two-derivative approximation. Clearly, if the entropy were simply the analog of the Bekenstein–Hawking area law we would obtain that all extended black-branes would have infinite entropy. Therefore, the way to obtain meaningful thermodynamics for such objects is to introduce regularized quantities in the form of worldvolume densities (see, for instance, deAntonioMartin:2012bi ; Meessen:2012su ). Reinstating Newton’s constant, we can conveniently define the entropy density $s$ for a $p$-brane living in $d$ spacetime dimensions as
@@
where $A_{\text{hor}}$ is the (dp2)(d-p-2)-dimensional black-brane transverse horizon, and $G_{d}$ is the $d$-dimensional Newton constant. Hence, for a five-dimensional black string one obtains
@@
where $\mathcal{R}_{\mu\nu}$ is the Ricci tensor. Equation (3.41) is not completely determined by the near-horizon geometry and to evaluate it we must know the precise structure of $\mathcal{L}_{3}$. Interestingly, this computation correctly reproduces the central charge (3.35) upon considering just the standard, two-derivative 5d $\mathcal{N}=1$ supergravity action supplemented with the known four-derivative corrections Castro:2007sd . Consequently, the macroscopic computation can be regarded as one-loop exact. On top of that, this confirms that (3.36) gives not only the leading term in the decompactification limit, but actually the exact entropy of a five-dimensional black string extended along an infinitely long compact circle of volume equal to $\mathcal{V}_{\rm h}$ in 5d Planck units.
@@ 3.4 Including non-perturbative effects
Our aim in the following will be to reconsider this point and show that, in fact, the main conclusions drawn from last section are left unchanged. To do so, we take two alternative routes that ultimately lead to the same answer. The first one proceeds, as explained in Section 3.4.1, by carefully evaluating the Schwinger determinant —in complex proper time— separately for each state in the D0-brane tower. Conversely, in Section 3.4.2, we derive an equivalent prescription by resumming the full tower of one-loop contributions, thereby treating both perturbative and non-perturbative corrections on an equal footing. We also briefly elaborate on the limitations associated to this second approach, which is very reminiscent of the recent proposal for computing the non-perturbative topological string partition function put forward in Hattab:2024ewk ; Hattab:2024chf ; Hattab:2024yol ; Hattab:2024ssg . For a lengthier discussion see Section 4.3.1 below.
@@ 3.4.1 Direct evaluation of the Schwinger integral
Our first strategy to obtain a non-perturbative definition of the leading-order prepotential at large volume proceeds similarly as we did in Section 3.3.1. There, following Gopakumar:1998ii ; Gopakumar:1998jq , we showed explicitly how by resorting to the dual M-theory description one is able to rewrite the relevant asymptotic series in an integral Schwinger-like form, cf. eqs. (3.25) and (3.26). Next, we performed the integration using an appropriate change of variable and a certain Fourier transform, which led us directly to the resummed perturbative expression (3.29). However, in doing so we were not concerned with some subtleties related to both the state-dependent change of integration variable (cf. footnote 3), as well as to possible singularities that could arise within the complex $s$-plane. In fact, it is easy to realize that, when relabeling $s\to s/2\pi n$ so as to reach the l.h.s. of (3.28), we must separate between states with $n\geq 0$ and $n<0$, since their contours cannot be simply deformed into one another due to the (double) poles arising from the hyperbolic sine in the denominator of the integrand. Hence, taking this into account leads to the following two distinct contributions within $\mathcal{I}(\alpha)$
@@
As a next step, and in order to be able to perform the Poisson resummation leading to the r.h.s. of (3.28), we need to deform the contour integral (3.42b) within the upper half plane so that it coincides with the positive real axis, thereby implying that we should pick up the residues of the infinitely many poles located $s=\frac{2\pi ik}{\alpha}$ for each $k\in\mathbb{N}$ (see Figure 4). The latter give rise to a non-perturbative contribution of the form [^19] Notice that (3.43) can be rewritten in terms of a single function $\varrho(\alpha)=i\frac{\alpha^{2}}{2\pi}\,\sum_{k=1}^{\infty}k^{-2}\left(1-e^{4\pi^{2}k/\alpha}\right)^{-1}$ as follows $\mathcal{I}^{(np)}(\alpha)=\alpha^{2}\frac{d}{d\alpha}\left(\frac{1}{\alpha}\,\varrho(\alpha)\right)\,.$
@@
For illustrative purposes, we have depicted the exact non-perturbative contribution, computed from (3.43), in Figure LABEL:sfig:ImaginarypartIcurve. Notice, in particular, the polynomial dependence with the expansion parameter $\alpha$ that arises in the deep five-dimensional regime, whose physical origin may seem surprising from the perspective of the auxiliary topological string theory. Furthermore, one might object that such divergent behavior could potentially undermine the discussion presented in Section 3.3.2, where it was argued that in the $\alpha\to\infty$ limit the entropy formula should match the one computed directly within the uplifted 5d supergravity theory. Very remarkably, we observe that this additional non-perturbative correction in $G(Y^{0},\Upsilon)$ does not modify the attractor equations nor the entropy function associated to the black hole system we are interested in here, since only the real part of $\mathcal{I}(\alpha)$ contributes to those, cf. eqs. (3.11)-(3.14). Let us stress here that the absence of this kind of corrections is intimately related to the non-perturbative stability of the BPS black hole background, a property that is ensured by supersymmetry. In any event, it is interesting to see explicitly how this expectation is borne out in the present set-up (see Section 4 for further evidence).
We also note, in passing, that the present analysis is in agreement with recent results obtained in Lin:2024jug , where it was shown that extremal Reissner-Nördstrom black holes exhibit a non-trivial spatial profile for their decay rate induced by non-perturbative emission of Swchinger pairs due to charged particles already existing in the theory, unless the latter are also extremal. Therefore, given that both the black hole solutions considered herein and the charged D0-branes fulfill this condition Gendler:2020dfp ; Heidenreich:2020upe ; Heidenreich:2024dmr , it makes perfect sense that such a decay channel does not exist in this case.